How can we convert from N/cm to MPa? or vice versa?

Can we convert for example 16 N/cm to MPa?
Answer #1

N/m2 are the SI unit of pressure.

other unit is pascal.

1 N/m2 = 1 pascal

And 1 m2 = 10 ^-4 cm2

Here we have to convert 1  N/cm2 = 10000 pacal

1 N/cm2 = 0.01 Mpascal

To convert 16 cm into Mpa

16 cm = 0.16 Mpa